OpenStudy (jenniferjuice):

(a − b)(a − b) is a^2 − b^2 always sometimes or never

OpenStudy (mertsj):

Did you do the multiplication and see what you get?

OpenStudy (anonymous):

(a − b)(a − b) a * a = a^2 a * -b = -ab -b * a = -ab -b * -b = + b so: a^2 - ab - ab + b then: a^2 - 2ab + b

OpenStudy (anonymous):

Do you see why?

OpenStudy (nory):

Two examples: (8-5)(8-5) what does that equal? (0-0) (0-0) and that?

OpenStudy (anonymous):

We have been doing these kind of problems, do you understand now??

OpenStudy (jenniferjuice):

so i wwould do the same i did for this one ike i did with the others? @some_someone

OpenStudy (anonymous):

yeah, basically you do multiplication, or FOIL

OpenStudy (jenniferjuice):

so it would be never?

OpenStudy (anonymous):

(a − b)(a − b) is a^2 − b^2 THIS IS FALSE.

OpenStudy (jenniferjuice):

yeh never...

OpenStudy (jenniferjuice):

well um thanks....:)

OpenStudy (jenniferjuice):

for your help